OpenStudy (anonymous):

5(b + 8) A. 5b + 40 B. 5b + 8 C. 5b + 13 D. 8b + 40

OpenStudy (uri):

Multiply 5 by the bracket so 5b +40.

OpenStudy (mayankdevnani):

\[\huge 5(b+8)\] \[\huge 5 \times b + 5 \times 8\] \[\huge 5b+40=answer\]

OpenStudy (goldphenoix):

Like uri said. You want to do the distributive property. So you distribute 5 to b and 8. 5 * b = 5b 5 * 8 = 40 Add them to get 5b+40

OpenStudy (mayankdevnani):

i hope you understand!! @flufffy95

OpenStudy (uri):

@mayankdevnani stop showing off.

OpenStudy (mayankdevnani):

what? @uri

OpenStudy (uri):

You know it.

OpenStudy (mayankdevnani):

i am helping that guy... i don't want to show off to anyone.. and also i don't like to show off

OpenStudy (uri):

Gooood for youuu.